# RouteKG: A knowledge graph-based framework for route prediction on road networks

### v3

@@ IV-B Data Preprocessing Module
To preserve the road network structure information, we construct the node adjacency edges (NAE) matrix, denoted as $\mathbf{A}$, which can be derived directly from the road network $\mathbf{G}$. To build $\mathbf{A}$, we pad the edges adjacent to each node to a uniform length, thereby creating a matrix of dimensions $\mathbb{R}^{|\mathbf{V}|\times N_{A}}$. In this context, $|\mathbf{V}|$ indicates the total number of nodes in the road network, while $N_{A}$ represents the maximum number of adjacent edges to any node. This padding approach enables batch training combined with smart masking techniques.
@@ IV-C Knowledge Graph Module
Intuitively, a driver’s route choice is based on their intended goal. Therefore, using KGC for route prediction aligns with the logic behind drivers’ route selections. However, most existing KGs are designed for search engines [73] and text-based Question Answering [74], making them unsuitable for direct application to road networks. Therefore, we need to construct a KG tailored to the characteristics of road networks, redefine the KGC problem in this context, and use learned spatial relations for more accurate route prediction. These tasks are encompassed in three submodules we’ve designed: Knowledge Graph Construction, Knowledge Graph Representation Learning, and Future Route Prediction through KGC. We will detail these in the following subsections.
@@ IV-C1 Knowledge Graph Construction
When constructing the KG $\mathcal{G}$ for road networks and routes, the initial key step is to identify entities $\mathcal{E}$. In the context of a road network, the predominant entity is the link. Every link $e$ is characterized by its unique identifiers and associated attributes such as length or connectivity. Selecting links as the sole entities reflects their intrinsic importance within the road network. It ensures the broad generalizability of the resulting KG across various routing contexts and scenarios, contributing to the applicability of the proposed approach.
@@
As discussed earlier, route prediction is reformulated as a KGC problem. Consequently, base on the selected entities (i.e., links), the relations $\mathcal{R}$ chosen for the road network should reflect and preserve the following features: (1) the spatial and structural properties of the road network, and (2) the consistency and preference patterns observed in drivers’ route selections. Given this, we identify four relations to construct the KG: connectivity, consistency, distance, and direction. Each relation offers unique insights into the relation between links within the road network. Connectivity describes the topological structure of links in the road network. A “ConnectBy” relation $\mathcal{R}_{c}$ is established between two links if they are directly connected via a shared node. The “ConsistentWith” relation $\mathcal{R}_{s}$ is derived from observed routes, capturing the co-occurrence of two links within the same routes. A higher co-occurrence rate indicates a stronger “ConsistentWith” relation, providing an empirical basis for capturing real-world routing patterns. The spatial distance between two links forms another key “DistanceTo” relation $\mathcal{R}_{a}$. It is crucial in many scenarios where the physical proximity of links impacts route selection and planning. Lastly, the direction forms the most critical “DirectionTo” relation $\mathcal{R}_{d}$ from a navigational standpoint. It provides essential directional information between links, greatly enhancing the graph’s utility for various routing tasks.
Table I summarizes the four major spatial relation types. By comprehensively capturing these four types of relations over the identified entities, the KG possesses a rich and nuanced representation of the road network, which could facilitate various advanced routing tasks.
@@
DistanceTo $\mathcal{R}^{a}$ Road Network $\mathbf{G}$
@@
TABLE I: Major spatial relation types and corresponding notations and data sources of the constructed KG. A major relation type may contain multiple relations. For example, the relation type “DirectionTo” contains $N_{d}$ directions, indicating a total of $N_{d}$ direction relations.
@@ IV-C2 Knowledge Graph Representation Learning
Spatial relations between entities (i.e. links) on a road network should be route-agnostic. This means that these relations should be independent of specific routes and instead solely reflect the spatial attributes of the road network itself. These relations also need to be encoded efficiently to support training and inference. One common approach is to employ KG embedding techniques, which aims to find embedding functions $\mathcal{M}_{\mathcal{E}}$, $\mathcal{M}_{\mathcal{R}}$ that maps each entity and each relation in to an feature vector. The embedding function $\mathcal{M}_{\mathcal{E}}(\cdot)$ and $\mathcal{M}_{\mathcal{E}}(\cdot)$ should preserve the inherent property of $\mathcal{G}$. However, road networks exhibit complex relations, as illustrated in Figure 3, involving many-to-many relations. To address this complexity, we modify and adapt the translation distance model TransH [52] to $\mathcal{G}$ in our study. This enables us to effectively learn the vector representations of both entities and relations in the knowledge graph $\mathcal{G}$.
Figure 3: An illustration of spatial relations, all of which are many-to-many.
To enable KG representation learning, we first need to construct the sets of positive triplets $\Delta$ and negative triplets $\Delta^{\prime}$ for each relation type. To facilitate batch representation learning and ensure comprehensive learning of all entities and relations, we employ a Random Sampling-based method for KG representation learning. The subsequent paragraphs will offer detailed descriptions of the positive and negative triplet sets’ construction processes for each relation type in $\mathcal{G}$.
@@
The ‘ConsistentWith’ relation $\mathcal{R}^{s}$, identifies co-occurring links in the same routes, indicating inter-links transition patterns. To construct $\Delta_{\mathcal{R}^{s}}$ and $\Delta^{\prime}_{\mathcal{R}^{s}}$, we utilize the spatial graph $\mathbf{G}$ and observed routes $\mathcal{X}^{o}$. The positive set $\Delta_{\mathcal{R}^{s}}$ consists of edges appearing together in observed routes, while the negative set $\Delta^{\prime}_{\mathcal{R}^{s}}$ is formed by randomly sampling edges from $\mathbf{G}$.
@@
For the ‘DistanceTo’ relation $\mathcal{R}^{a}$, sampling is done from observed routes $\mathcal{X}^{o}$, better aligning the “DistanceTo” relation with route prediction and reducing the possible $\mathcal{R}^{a}$ relations. The resulting sets are $\Delta_{\mathcal{R}^{a}}$ and $\Delta^{\prime}_{\mathcal{R}^{a}}$ for positive and negative triplets, respectively.
@@
We need the inter-link direction matrix $\mathbf{D}$ to construct positive and negative triplet sets $\Delta_{\mathcal{R}^{d}}$ and $\Delta^{\prime}_{\mathcal{R}^{d}}$. For sampled edges $e^{i},e^{j}\in\mathbf{E}$, their relative direction $r^{d}$ is given by $r^{d}=\mathbf{D}_{e^{i}e^{j}}$, forming the positive set. The negative set is similarly formed from edge pairs that contradict the directional relation in $\mathbf{D}$.
We denote the positive triplet sets for all relations as $\Delta_{\mathcal{R}^{\cdot}}$ and the negative sets as $\Delta_{\mathcal{R}^{\cdot}}^{\prime}$. Consider an identified type of relation $\mathcal{R}^{\cdot}$, and we incorporate two trainable weight matrices. One matrix functions as the relation embedding matrix, represented as $\mathbf{W}_{\mathcal{R}^{\cdot}}\in\mathbb{R}^{|\mathcal{R}^{\cdot}|\times\delta_{\mathcal{R}^{\cdot}}}$, while the other corresponds to the relation hyperplane, denoted as $\mathbf{P}_{\mathcal{R}^{\cdot}}$, both maintaining congruent dimensions. Given the sets of positive triplets $\Delta_{\cdot}$ and negative triplets $\Delta_{\cdot}^{\prime}$, where $\cdot$ can represent any of the relations on the KG, the representation learning process involves the following steps. For any triplet $(h,r,t)\in\Delta_{\cdot}$ and $(h^{\prime},r^{\prime},t^{\prime})\in\Delta^{\prime}_{\cdot}$, we use $\mathbf{h}$, $\mathbf{r}$, $\mathbf{t}$, $\mathbf{h}^{\prime}$, $\mathbf{r}^{\prime}$, and $\mathbf{t}^{\prime}$ denoted their embeddings and use $\mathbf{p}^{r}$ and $\mathbf{p}^{r^{\prime}}$ to denote the hyperplane of relation $r$ and $r^{\prime}$ with $\left(\mathbf{p}^{r}\right)^{\top}$ and $\left(\mathbf{p}^{r^{\prime}}\right)^{\top}$ as their transposes. Then we carry out the process defined by Eq. (4) for all relations present on the KG. The loss function for the KG $\mathcal{G}$’s representation learning is:
@@
[ $\left(\mathbf{h}-\left(\mathbf{p}^{r}\right)^{\top}\mathbf{h}\mathbf{p}^{r}\right)+\mathbf{r}-\left(\mathbf{t}-\left(\mathbf{p}^{r}\right)^{\top}\mathbf{t}\mathbf{p}^{r}\right)\text{ }\right\|_{\ell_{1}}+\psi-$
 $\left(\mathbf{h}^{\prime}-(\mathbf{p}^{r^{\prime}})^{\top}\mathbf{h}^{\prime}\mathbf{p}^{r^{\prime}}\right)+\mathbf{r}^{\prime}-\left(\mathbf{t}^{\prime}-(\mathbf{p}^{r^{\prime}})^{\top}\mathbf{t}^{\prime}\mathbf{p}^{r^{\prime}}\right)\text{ }\right\|_{\ell_{1}}\text{ }\Biggr{]}_{+},$
@@ IV-C3 Future Route Prediction through Knowledge Graph Completion
Our study approaches the task of future route prediction by framing it as a KGC problem. As shown in Figure 4, given the last link $e_{i}^{\Gamma}$ (i.e., head entity) of the $i$-th observed route $x^{o}_{i}$ and the (estimated or actual) direction of movement (i.e., relation), our objective is to infer the future route $\widehat{x^{f}_{i}}$ (i.e., tail entity) that the vehicle will traverse. In our case, the actual direction $r^{d}_{i}$ of a vehicle’s movement is the direction from the current link to the last link of the future route. Utilizing KGC, we introduce an innovative objective to predict the immediate future routes of road users. This addition not only utilizes the learned KG embeddings to enhance route prediction accuracy but also enriches the KG representation with deeper semantics, thereby creating a synergistic effect between KG embedding and route prediction.
@@
For illustration, we consider the $i$-th observed route $x^{o}_{i}\in\mathcal{X}^{o}$ and the corresponding direction $x^{o,d}_{i}\in\mathcal{X}^{o,d}$, drawn from the set of observed routes $\mathcal{X}^{o}$ and directions $\mathcal{X}^{o,d}$, respectively. Here, $x^{o}_{i}=\{e^{j}_{i}\}^{\Gamma}_{j=1}$ and $x^{o,d}_{i}=\{e^{d,j}_{i}\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}$. Initially, we extract the embeddings of all elements of $x^{o}$ and $x^{o,d}$ by multiplying their respective one-hot vectors with the corresponding trainable embedding matrices: $\mathbf{W}_{\mathcal{E}}\in\mathbb{R}^{|\mathcal{E}|\times\delta_{\mathcal{E}}}$ and $\mathbf{W}_{\mathcal{R}^{d}}\in\mathbb{R}^{|\mathcal{R}^{d}|\times\delta_{\mathcal{R}^{d}}}$. The resulting embeddings for $x^{o}_{i}$ and $x^{o,d}_{i}$ are denoted as $\mathbf{x}^{o}_{i}=\{\mathbf{e}^{j}_{i}\|\}_{j=1}^{\Gamma}$ and $\mathbf{x}^{o,d}_{i}=\{\mathbf{e}^{d,j}_{i}\|\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}$, with $\mathbf{e}_{i}^{\cdot}\in\mathbb{R}^{\delta_{\mathcal{E}}}$, $\mathbf{e}^{d,\cdot}_{i}\in\mathbb{R}^{\delta_{\mathcal{R}^{d}}}$. To predict the direction of the vehicle’s future routes, we utilize a Multi-layer Perceptron (MLP) [75] to encode $\mathbf{x}^{o}_{i}$ and $\mathbf{x}^{o,d}_{i}$:
@@
where $\widehat{r^{d}_{i}}$ represents the estimated direction of the $i$-th future route, and we employ the cross-entropy loss to optimize the parameters of the $\mathrm{MLP}_{d}$:
@@
It should be noted that the estimation of the vehicle’s future route direction is only necessary when the goal is unspecified (i.e., subproblem 1). Conversely, when the goal direction or the actual goal is provided, one can directly utilize the given goal direction (i.e., subproblems 2 and 3).
To employ KGC to predict future routes, the last link $e_{i}^{\Gamma}$ is converted into the corresponding entity embedding by multiplying the one-hot vector of $e_{i}^{\Gamma}$ with $\mathbf{W}_{\mathcal{E}}$, yielding $\mathbf{e}_{i}^{\Gamma}\in\mathbb{R}^{\delta_{\mathcal{E}}}$. Then, $\mathbf{r}^{d}_{i}\in\mathbb{R}^{\delta_{\mathcal{R}^{d}}}$ and $\mathbf{p}^{d}_{i}\in\mathbb{R}^{\delta_{\mathcal{R}^{d}}}$ are obtained through a similar operation with $\mathbf{W}_{\mathcal{R}^{d}}$ and $\mathbf{P}_{\mathcal{R}^{d}}$, respectively. Note that in this stage, RouteKG only needs the current position of a vehicle (i.e., the last link of the observed route), which is fundamentally different from existing seq2seq methods. Given $\mathbf{p}^{d}_{i}$, we first project the $\mathbf{e}_{i}^{\Gamma}\in\mathbb{R}^{\delta_{\mathcal{E}}}$ and all links embeddings $\mathbf{e}\in\mathbb{R}^{|\mathcal{E}|\times\delta_{\mathcal{E}}}$ to the hyperplane $\mathbf{p}^{d}_{i}$ to obtain the projected head embedding $\mathbf{e}_{i,\perp}^{\Gamma}$ and all candidate tail embedding $\mathbf{e}_{\perp}$ on the hyperplane:
@@
where $\mathbf{e}_{i,\perp}^{\Gamma}\in\mathbb{R}^{\delta_{\mathcal{E}}}$ and $\mathbf{e}_{\perp}\in\mathbb{R}^{|\mathcal{E}|\times\delta_{\mathcal{E}}}$.
Upon acquiring the projected head embedding $\mathbf{e}_{i,\perp}^{\Gamma}$, we add the relation to the projected head embedding to query the tail entity. Given the projected head embedding $\mathbf{e}_{i,\perp}^{\Gamma}$, the direction relation embedding $\mathbf{r}^{d}_{i}$, and the distance relation embedding $\mathbf{r}^{a,\gamma}_{i}$, we could query the tail entity based on the following equation:
$\mathrm{Pr}(\widetilde{x^{f,\gamma}_{i}})=\mathrm{Softmax}\left(\mathbf{e}_{\perp}\cdot\left[\left(\mathbf{e}_{i,\perp}^{\Gamma}+\mathbf{r}^{d}_{i}\right)\odot\mathbf{r}^{a,\gamma}_{i}\right]^{\top}\right),$ (8)
where $\mathrm{Pr}(\widetilde{x^{f,\gamma}_{i}})\in\mathbb{R}^{|\mathcal{E}|}$ is the predicted probability distribution which indicates the likelihood of each link being the $\gamma$-th link of the $i$-th future route. We can set $\gamma$ from 1 to $\Gamma^{\prime}$ and recursively use Eq. (6) to obtain $\Gamma^{\prime}$ probability distributions $\{\mathrm{Pr}(\widetilde{x^{f,\gamma}_{i}})\}_{\gamma=1}^{\Gamma^{\prime}}$ representing the estimated future route probabilities, which is the final output of the module.
@@
where $x_{i}^{f,\gamma}$ is the actual $\gamma$-th link of the $i$-th future route, and []delimited-[][$\cdot$][ ⋅ ] is the indexing operation that retrieves and maximizes the log-probability of the actual link. Note that $x_{i}^{f,\gamma}$ and $e_{i}^{\Gamma+\gamma}$ indicate the same link in the $i$-th future route.

### v4

@@ IV-B Data Preprocessing Module
To preserve the road network structure information, we construct the node adjacency edges (NAE) matrix, denoted as $\mathbf{A}$, which can be derived directly from the road network $\mathbf{G}$. To build $\mathbf{A}$, we pad the edges adjacent to each node to a uniform length, thereby creating a matrix of dimensions $\mathbb{R}^{|\mathbf{V}|\times N_{A}}$. In this context, $|\mathbf{V}|$ indicates the total number of nodes in the road network, while $N_{A}$ represents the maximum number of adjacent edges to any node. This padding approach enables batch training combined with smart masking techniques. This matrix not only encodes structural adjacency but also enables efficient batch retrieval of neighboring edge attributes (including embeddings), which is critical for learning relation-aware representations and route reranking.
@@ IV-C Knowledge Graph Module
Intuitively, a user’s route choice is based on their intended goal. Therefore, using KGC for route prediction aligns with the logic behind drivers’ route selections. However, most existing KGs are designed for search engines [71] and text-based Question Answering [72], making them unsuitable for direct application to road networks. Therefore, we need to construct a KG tailored to the characteristics of road networks, redefine the KGC problem in this context, and use learned spatial relations for more accurate route prediction. These tasks are encompassed in three submodules we’ve designed: Knowledge Graph Construction, Knowledge Graph Representation Learning, and Future Route Prediction through KGC. We will detail these in the following subsections.
@@ IV-C1 Knowledge Graph Construction
When constructing the KG $\mathcal{G}$ for road networks and routes, the initial step is to identify entities $\mathcal{E}$. In the context of a road network, the predominant entity is the link. Each link $e$ is characterized by its unique identifiers and associated attributes such as length or connectivity. Selecting links as the sole entities reflects their intrinsic importance within the road network. It also ensures the generalizability of the proposed approach in various contexts and scenarios.
@@
As discussed earlier, route prediction is reformulated as a KGC problem. Based on the selected entities (i.e., links), the relations $\mathcal{R}$ chosen for the road network should reflect and preserve the following features: (1) the spatial and structural properties of the road network, and (2) the consistency and preference patterns in observed route selections. Given this, we identify four relations to construct the KG, as illustrated in Table I.
TABLE I: Major spatial relation types, corresponding notations and data sources.
@@
DistanceTo $\mathcal{R}^{a}$ Road Network $\mathbf{G}$, Observed routes $\mathcal{X}^{o}$
@@
The ConnectBy relation, denoted as $\mathcal{R}_{c}$, defines whether two links are directly connected via a shared node (intersection). Mathematically, for links $e^{i}$ and $e^{j}$, this relation is represented as: $(e^{i},\mathcal{R}_{c},e^{j})=\begin{cases}1,&\text{if }e^{i}\text{ and }e^{j}\text{ share a node}\\ 0,&\text{otherwise}\end{cases}$. This relation models the physical connectivity between links, helping the model understand the road network’s layout.
The ConsistentWith relation, denoted as $\mathcal{R}_{s}$, captures the co-occurrence of two links in the same observed routes. A higher co-occurrence rate indicates a stronger ConsistentWith relation. This relation is defined as: $(e^{i},\mathcal{R}_{s},e^{j})=\frac{|\{x\in\mathcal{X}^{o}:e^{i}\in x,e^{j}\in x\}|}{|\mathcal{X}^{o}|},$ where $\mathcal{X}^{o}$ is the set of observed routes. This reflects driver behavior, indicating which links are commonly used together in routes, and helps capture real-world routing patterns.
The DistanceTo relation, denoted as $\mathcal{R}_{a}$, measures the physical distance between two links. For two links $e^{i}$ and $e^{j}$, it is defined as the Euclidean or shortest-path distance: $(e^{i},\mathcal{R}_{a},e^{j})=d(e^{i},e^{j}),$ where $d(\cdot,\cdot)$ represents a distance metric, which can be physical distance or network distance. This relation reflects the spatial proximity of links and helps the model understand how distance affects route selection.
The DirectionTo relation, denoted as $\mathcal{R}_{d}$, represents the navigational direction between two links. For two links $e^{i}$ and $e^{j}$, it is defined as a discrete direction class: $(e^{i},\mathcal{R}_{d},e^{j})=dir(e^{i},e^{j}),$ where $dir(e^{i},e^{j})$ represents the discretized direction class. This relation encodes navigational decisions, such as turns or straight paths, which are crucial for route planning.
Note that each relation type may contain multiple relations. For example, the relation type “DirectionTo” contains $N_{d}$ directions, indicating a total of $N_{d}$ direction relations. By comprehensively capturing these four types of relations, the KG offers a rich and nuanced representation of the road network, which can facilitate various routing tasks.
@@ IV-C2 Knowledge Graph Representation Learning
Spatial relations between entities (i.e. links) on a road network should be route-agnostic. This means that these relations should be independent of specific routes and instead solely reflect the spatial attributes of the road network itself. These relations also need to be encoded efficiently to support training and inference. One common approach is to employ KG embedding techniques, which aim to find embedding functions $\mathcal{M}_{\mathcal{E}}$, $\mathcal{M}_{\mathcal{R}}$ that map each entity and each relation into a feature vector. The embedding function $\mathcal{M}_{\mathcal{E}}(\cdot)$ and $\mathcal{M}_{\mathcal{R}}(\cdot)$ should preserve the inherent property of $\mathcal{G}$. However, road networks exhibit complex relations, as illustrated in Figure 3, involving many-to-many relations. To address this complexity, we modify and adapt the translation distance model TransH [50] in our study, so that we can effectively learn the vector representations of both entities and relations in $\mathcal{G}$.
Figure 3: An illustration of many-to-many spatial relations between sampled head and tail entities. Each relation captures specific patterns between entity pairs: (a) ConnectBy: adjacency in the road network via a shared node; (b) ConsistentWith: co-occurrence in observed routes; (c) DistanceTo: spatial or network distance; (d) DirectionTo: navigational direction.
To enable KG representation learning, we first need to construct sets of positive triplets $\Delta$ and negative triplets $\Delta^{\prime}$ for each relation type. Positive triplets reflect connections (e.g., physically connected links or frequently co-used links), while negative triplets are generated by sampling incorrect relationships. To facilitate batch representation learning and ensure comprehensive learning of all entities and relations, we employ a random sampling-based method. The specific construction process for each relation type is described below.
@@
$\mathcal{R}^{s}$ captures transition patterns between links that frequently co-occur within the same observed route, indicating a form of spatial or behavioral consistency. To construct the positive set $\Delta_{\mathcal{R}^{s}}$ and negative set $\Delta^{\prime}_{\mathcal{R}^{s}}$, we utilize the spatial graph $\mathbf{G}$ and observed routes $\mathcal{X}^{o}$. The positive set $\Delta_{\mathcal{R}^{s}}$ contains pairs of links that co-occur in the same observed route. To construct the negative set $\Delta^{\prime}_{\mathcal{R}^{s}}$, we randomly sample link pairs from $\mathbf{G}$ that do not co-occur in any observed route in $\mathcal{X}^{o}$. While we do not explicitly encode co-occurrence frequency, frequent entity pairs are more likely to be sampled, which implicitly reflects the defined $\mathcal{R}^{s}$ relation.
@@
For $\mathcal{R}^{a}$, sampling is performed using observed routes $\mathcal{X}^{o}$, better aligning the “DistanceTo” relation with route prediction and reducing the possible $\mathcal{R}^{a}$ relations. The resulting sets are $\Delta_{\mathcal{R}^{a}}$ and $\Delta^{\prime}_{\mathcal{R}^{a}}$ for positive and negative triplets, respectively.
@@
We need the inter-link direction matrix $\mathbf{D}$ to construct positive and negative triplet sets $\Delta_{\mathcal{R}^{d}}$ and $\Delta^{\prime}_{\mathcal{R}^{d}}$. For sampled edges $e^{i},e^{j}\in\mathbf{E}$, their relative direction $r^{d}$ is given by $r^{d}=\mathbf{D}_{e^{i}e^{j}}$, forming the positive set. The negative set is formed using edge pairs that contradict the directional relation in $\mathbf{D}$ (i.e., opposite direction).
We denote the positive triplet sets for all relations as $\Delta_{\mathcal{R}^{\cdot}}$ and the negative sets as $\Delta_{\mathcal{R}^{\cdot}}^{\prime}$. Consider an identified type of relation $\mathcal{R}^{\cdot}$, and we incorporate two trainable weight matrices. One matrix functions as the relation embedding matrix, represented as $\mathbf{W}_{\mathcal{R}^{\cdot}}\in\mathbb{R}^{|\mathcal{R}^{\cdot}|\times\delta_{\mathcal{R}^{\cdot}}}$, while the other corresponds to the relation hyperplane, denoted as $\mathbf{P}_{\mathcal{R}^{\cdot}}$, both maintaining congruent dimensions. The relation hyperplane $\mathbf{P}_{\mathcal{R}^{\cdot}}$ defines a learned subspace onto which entity embeddings are projected, enabling the model to capture relational constraints in a more structured and discriminative manner. Given the sets of positive triplets $\Delta_{\cdot}$ and negative triplets $\Delta_{\cdot}^{\prime}$, where $\cdot$ can represent any of the relations on the KG, the representation learning process involves the following steps. For any triplet $(h,r,t)\in\Delta_{\cdot}$ and $(h^{\prime},r^{\prime},t^{\prime})\in\Delta^{\prime}_{\cdot}$, we use $\mathbf{h}$, $\mathbf{r}$, $\mathbf{t}$, $\mathbf{h}^{\prime}$, $\mathbf{r}^{\prime}$, and $\mathbf{t}^{\prime}$ denoted their embeddings and use $\mathbf{p}^{r}$ and $\mathbf{p}^{r^{\prime}}$ to denote the hyperplane of relation $r$ and $r^{\prime}$ with $\left(\mathbf{p}^{r}\right)^{\top}$ and $\left(\mathbf{p}^{r^{\prime}}\right)^{\top}$ as their transposes. For all relations, the loss function for learning the representation of the KG $\mathcal{G}$ is:
@@
[ $\Biggl[\left\|\text{ }\left(\mathbf{h}-\left(\mathbf{p}^{r}\right)^{\top}\mathbf{h}\mathbf{p}^{r}\right)+\mathbf{r}-\left(\mathbf{t}-\left(\mathbf{p}^{r}\right)^{\top}\mathbf{t}\mathbf{p}^{r}\right)\text{ }\right\|_{\ell_{1}}+\psi$-
[…]
@@ IV-C3 Future Route Prediction through Knowledge Graph Completion
Our study approaches the task of future route prediction by framing it as a KGC problem. As shown in Figure 4, given the last link $e_{i}^{\Gamma}$ (i.e., head entity) of the $i$-th observed route $x^{o}_{i}$ and the (estimated or actual) direction of movement (i.e., relation), our objective is to infer the future route $\widehat{x^{f}_{i}}$ (i.e., tail entity) that the user will traverse. In our case, the actual direction $r^{d}_{i}$ of a user’s movement is the direction from the current link to the last link of the future route. Utilizing KGC, we introduce an innovative objective to predict the immediate future routes of road users. This addition not only utilizes the learned KG embeddings to enhance route prediction accuracy but also enriches the KG representation with deeper semantics, thereby creating a synergistic effect between KG embedding and route prediction.
@@
We use the EGoalD for the whole process illustration and then provide the details for the GoalD and Goal settings.
EGoalD Setting
We consider the $i$-th observed route $x^{o}_{i}\in\mathcal{X}^{o}$ and its corresponding direction sequence $x^{o,d}_{i}=\left\{e^{d,j}_{i}\right\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}\in\mathcal{X}^{o,d}$, where $x^{o}_{i}=\left\{e^{j}_{i}\right\}^{\Gamma}_{j=1}$ represents a sequence of links and $e^{d,j}_{i}$ denotes the direction label of the $j$-th link. Here, $\mathcal{X}^{o,d}$ denotes the set of all such direction sequences. Initially, we extract the embeddings of all elements of $x^{o}$ and $x^{o,d}$ by multiplying their respective one-hot vectors with the corresponding trainable embedding matrices: $\mathbf{W}_{\mathcal{E}}\in\mathbb{R}^{|\mathcal{E}|\times\delta_{\mathcal{E}}}$ and $\mathbf{W}_{\mathcal{R}^{d}}\in\mathbb{R}^{|\mathcal{R}^{d}|\times\delta_{\mathcal{R}^{d}}}$. The resulting embeddings for $x^{o}_{i}$ and $x^{o,d}_{i}$ are denoted as $\mathbf{x}^{o}_{i}=\{\mathbf{e}^{j}_{i}\|\}_{j=1}^{\Gamma}$ and $\mathbf{x}^{o,d}_{i}=\{\mathbf{e}^{d,j}_{i}\|\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}$, with $\mathbf{e}_{i}^{\cdot}\in\mathbb{R}^{\delta_{\mathcal{E}}}$, $\mathbf{e}^{d,\cdot}_{i}\in\mathbb{R}^{\delta_{\mathcal{R}^{d}}}$.
To predict the direction of the user’s future routes, we utilize a Multi-layer Perceptron (MLP) [73] to encode $\mathbf{x}^{o}_{i}$ and $\mathbf{x}^{o,d}_{i}$:
@@
where $\|$ is the concatenate operation, $\widehat{r^{d}_{i}}$ represents the estimated direction of the $i$-th future route, and we employ the cross-entropy loss to optimize the parameters of the $\mathrm{MLP}_{d}$:
@@
To predict future routes, the last link $e_{i}^{\Gamma}$ is converted into the corresponding entity embedding by multiplying the one-hot vector of $e_{i}^{\Gamma}$ with $\mathbf{W}_{\mathcal{E}}$, yielding $\mathbf{e}_{i}^{\Gamma}\in\mathbb{R}^{\delta_{\mathcal{E}}}$. Then, $\mathbf{r}^{d}_{i}\in\mathbb{R}^{\delta_{\mathcal{R}^{d}}}$ and $\mathbf{p}^{d}_{i}\in\mathbb{R}^{\delta_{\mathcal{R}^{d}}}$ are obtained through a similar operation with $\mathbf{W}_{\mathcal{R}^{d}}$ and $\mathbf{P}_{\mathcal{R}^{d}}$, respectively. Note that in this stage, RouteKG only needs the user’s current position (i.e., the last link of the observed route), which is fundamentally different from existing seq2seq methods. Given $\mathbf{p}^{d}_{i}$, we first project the $\mathbf{e}_{i}^{\Gamma}\in\mathbb{R}^{\delta_{\mathcal{E}}}$ and all links embeddings $\mathbf{e}\in\mathbb{R}^{|\mathcal{E}|\times\delta_{\mathcal{E}}}$ to the hyperplane $\mathbf{p}^{d}_{i}$ to obtain the projected head embedding $\mathbf{e}_{i,\perp}^{\Gamma}$ and all candidate tail embedding $\mathbf{e}_{\perp}$ on the hyperplane:
@@
where $\mathbf{e}_{i,\perp}^{\Gamma}\in\mathbb{R}^{\delta_{\mathcal{E}}}$ and $\mathbf{e}_{\perp}\in\mathbb{R}^{|\mathcal{E}|\times\delta_{\mathcal{E}}}$. The projection ensures that the similarity computation is constrained to a direction-specific subspace, allowing the model to capture semantics aligned with the intended travel direction.
Upon acquiring the projected head embedding $\mathbf{e}_{i,\perp}^{\Gamma}$, we add the relation to the projected head embedding to query the tail entity. Given the projected head embedding $\mathbf{e}_{i,\perp}^{\Gamma}$, the direction relation embedding of the estimated direction $\widehat{\mathbf{r}^{d}_{i}}$, and the distance relation embedding $\mathbf{r}^{a,\gamma}_{i}$, we could query the tail entity based on the following equation:
$\mathrm{Pr}(\widetilde{x^{f,\gamma}_{i}})=\mathrm{Softmax}\left(\mathbf{e}_{\perp}\cdot\left[\left(\mathbf{e}_{i,\perp}^{\Gamma}+\widehat{\mathbf{r}^{d}_{i}}\right)\odot\mathbf{r}^{a,\gamma}_{i}\right]^{\top}\right),$ (8)
[…]
where $x_{i}^{f,\gamma}$ is the actual $\gamma$-th link of the $i$-th future route, and the indexing operation $\cdot$] retrieves the predicted log-probability assigned to the ground-truth link, and the loss corresponds to maximizing the likelihood of the correct link via a standard cross-entropy objective. Note that $x_{i}^{f,\gamma}$ and $e_{i}^{\Gamma+\gamma}$ indicate the same link in the $i$-th future route.
GoalD Setting
It should be noted that the estimation of the user’s future route direction is only necessary when the goal is unspecified (i.e., subproblem 1). Conversely, when the goal direction or the actual goal is provided, one can directly utilize the given goal direction (i.e., subproblems 2 and 3). The embedding of the estimated direction $\widehat{\mathbf{r}^{d}_{i}}$ in Eq. (9) can be replaced by the actual direction relation embedding $\mathbf{r}^{d}_{i}$.
Goal Setting